# Magic state distillation without measurements and post-selection

### v1

@@ Abstract
Clifford gates are assumed to be noise-free as the result of employing MSD on logical qubits that are encoded in a QEC code [^2] Note that this logical qubit QEC code is, in general, different from the distillation QEC code. with transversal Clifford gates in the limit $d\rightarrow\infty$.
@@
We anticipate that the unitary decoding circuit could be optimized with regards to the number of CNOT gates and circuit depth because, for our application, only the four $X$-syndromes are required to correct $Z$-flips. Moreover, the design of the coherent feedback network has been performed manually for this work. This process could be automated by classical optimization techniques and might then easily be adapted to other QEC codes [66]. It would be interesting to investigate the scheme on other MSD protocols, especially with more recently discovered QEC code constructions for transversal non-Clifford gates such as [25, 67].
As our coherent feedback operations are embodied by mostly multi-qubit-controlled gates, such a construction seems convenient for application on logical qubits that are encoded in a QEC code with a suitable fault-tolerant construction of these gates; see for instance the construction of FT $C^{k}Z$ gates in the Bacon-Shor code given in Ref. [17]. Although these gates are not simply transversal, the authors suggest fault-tolerant constructions by consecutively extending and shrinking code patches for arbitrary code distances.
@@ Acknowledgements
Appendix A Physical circuit for measurement-free 15-to-1 magic state distillation
@@
Figure 3: Explicit circuit to perform 15-to-1 MSD (see Fig. 1) via the inverse unitary encoding block $E_{15}^{\dagger}$ followed by the coherent feedback network (CFN). The CFN serves to deterministically correct every Pauli error on the output qubit 0, which may result from an arbitrary single-qubit error on any of the 15 input qubits. Qubits 1 to 14 are in the state |$\ket{0}$ after $E_{15}^{\dagger}$ in the absence of noise and are discarded at the end of the circuit. As an example, the input error $Z_{1}$ (blue star) is shown to propagate to qubit 0 while flipping qubits 1 and 8 (see fifth row of Tab. 1) and triggering a combination of three multi-qubit-controlled gates in the CFN (red circles) that corrects the error on the output qubit 0. The multi-qubit-controlled-$X$ gates can be left out when dealing with the noise model of Eq. (1).
@@
Table 1: For the 15 possible single-qubit input errors of $X$- and $Z$-type, only a subset of errors flip the output message qubit 0 while also flipping some of the syndrome qubits 1 to 14. The coherent feedback network in Fig. 3 implements these syndrome-conditioned flips while respecting the "no flip" conditions.

### v2

@@ Abstract
Such a CFN seems convenient for application on logical qubits that are encoded in a QEC code with a suitable FT construction of multi-qubit-controlled gates; see for instance the construction of FT $C^{k}Z$ gates in the Bacon-Shor code given in Refs. [59, 17]. Although these gates are not simply transversal, the authors suggest FT constructions by consecutively extending and shrinking code patches for arbitrary code distances. In Fig. 4, we additionally provide a simplified implementation based on coherently-controlled-reset operations that does not use multi-qubit-controlled gates.
@@
Gates are assumed to be noise-free as the result of employing MSD on logical qubits that are encoded in an $[[N,K,D]]$ QEC code [^2] Note that this logical qubit QEC code is, in general, different from the distillation QEC code. in the limit $D\rightarrow\infty$.
@@
The design of the coherent feedback network has been performed manually for this work. We anticipate that this process could be automated by classical optimization techniques and might then easily be adapted to other QEC codes [67]. It would be interesting to investigate the scheme on other MSD protocols, especially with more recently discovered code constructions for transversal non-Clifford gates such as [25, 68]. Our concept of unitary decoding with subsequent message qubit corrections could be translated to an entanglement distillation scenario [69, 70, 71].
@@ Acknowledgements
Appendix A Physical circuits for measurement-free 15-to-1 magic state distillation
@@
Figure 3: Explicit circuit to perform 15-to-1 MSD (see Fig. 1) via the inverse unitary encoding block $E_{15}^{\dagger}$ followed by the coherent feedback network (CFN). The CFN serves to deterministically correct every Pauli error on the output qubit 0, which may result from an arbitrary single-qubit error on any of the 15 input qubits. Qubits 1 to 14 are in the state |$\ket{0}$ after $E_{15}^{\dagger}$ in the absence of noise and are discarded at the end of the circuit. As an example, the input error $Z_{1}$ (blue star) is shown to propagate to qubit 0 while flipping qubits 1 and 8 (see fifth row of Tab. 1) and triggering a combination of three multi-qubit-controlled gates in the CFN (red circles) that corrects the error on the output qubit 0. The multi-qubit-controlled-$X$ gates can be left out when dealing with the noise model of Eq. (1). The CNOT gates marked gray do not actually propagate logical operations. They can be left out without compromising the ability to distinguish single-qubit $Z$-errors (see Fig. 4).
@@
Table 1: For the 15 possible single-qubit input errors of $X$- and $Z$-type, only a subset of errors flip the output message qubit 0 while also flipping some of the syndrome qubits 1 to 14. The coherent feedback network in Fig. 3 implements these syndrome-conditioned flips while respecting the "no flip" conditions. No feedback can be applied when all syndrome qubits are in the |$\ket{0}$ state.
Figure 4: Simplified distillation circuit to correct $Z$-errors only. Unnecessary CNOT gates from $E_{15}^{\dagger}$ in Fig. 3 are removed and the resulting circuit is called $D_{15}^{\dagger}$. The CFN can be implemented without multi-qubit-controlled gates if a coherently-controlled-reset ($R$) operation is available. Conditioned on qubit 14, the other three syndrome qubits 1, 2 and 8 are reset. The first controlled-$Z$ operation therefore only corrects qubit 0 in case that syndrome qubits $1,\{1,2\},\{1,8\}$ or $\{1,2,8\}$ are flipped and then qubits 2 and 8 are reset so that no further feedback is applied. If qubit 1 is in the state |$\ket{0}$, the second controlled-$Z$ operation is only applied when the syndrome qubits $2$ or $\{2,8\}$ are flipped and qubit 8 is reset. If also qubit 2 is in the state |$\ket{0}$, the last controlled-$Z$ operation only acts if the syndrome qubit $8$ is flipped.
We note that some CNOT gates in Fig. 3, while mapping stabilizers back to their corresponding physical qubits, need not actually be considered in the context of decoding because they do not propagate physical $Z$-errors to the message qubit 0. Removing these CNOT leaves us with the same correction conditions for the CFN as before: Apply a $Z$-flip to qubit 0 if qubit 1 OR 2 OR 8 are flipped AND qubit 14 is not flipped. If qubit 14 is flipped OR no $X$-syndrome qubits are flipped, apply no $Z$-flip to qubit 0. In Fig. 4, we avoid usage of multi-qubit-controlled gates and instead show an implementation of this CFN by means of a coherently-controlled-reset operation.